OpenStudy (anonymous):

solve 8x-(5x+8)=4. i came up with x=4?

OpenStudy (anonymous):

yup. correct! i checked it for you.

OpenStudy (anonymous):

8x-5x-8=4 3x = 4+8 x =12/3 x =4

OpenStudy (anonymous):

yay thanks!

OpenStudy (anonymous):

np